OpenStudy (anonymous):

Write the quadratic function in vertex form. y = x^2 - 6x + 7

OpenStudy (anonymous):

Use completion of square method: \[y = x^2 - 6x + 3^2 - 3^2 + 7\] \[y = (x-3)^2 - 2\]

OpenStudy (anonymous):

thats all u do?

OpenStudy (anonymous):

y=x^2-6x+7

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

again the same formula to find the x coordinate of the vertex just use the formula...(i think u know it by now!)  then, x =  plug that value in the original equation to find y coordinate! y = 3^2 -6(3) + 7... equals -2!